OpenStudy (anonymous):

which of the following is the solution to the inequality x^2+6x≥7?

OpenStudy (anonymous):

folowing?

OpenStudy (anonymous):

sorry wait a minute what is the the solution to the inequality

OpenStudy (anonymous):

math book im new

OpenStudy (anonymous):

do u hve the answers,

OpenStudy (anonymous):

yeah a few

OpenStudy (anonymous):

show them, here...

OpenStudy (anonymous):

x≥7 or x≤-1 x≤-7 or x≥1 x≤7 and x≥-1 x≥-7 and x≤1

OpenStudy (anonymous):

@satellite73 ...Check if i am write \[x^{2}+6x \ge7\] \[x(x+6)\ge 7\] \[x \ge 7 or x \ge 1\]

OpenStudy (anonymous):

sorry no not right

OpenStudy (anonymous):

of course my answer was wrong too should be \[x^2+6x-7\geq0\] \[(x+7)(x-1)\geq 0\] \(x\leq -7\) or \(x\geq 1\)

OpenStudy (anonymous):

how can the sign change...?

OpenStudy (saifoo.khan):

Sat got wrong? D: is it Armageddon?

OpenStudy (anonymous):

zeros of the quadratic are -7 and 1 and we know that the graph is a parabola that faces up and looks something like this 

OpenStudy (anonymous):

so it is negative between the zeros i.e. for \(-7<x<1\) and positive outside them, i.e. for \(x\leq-7\) or \(x\geq 1\)

OpenStudy (anonymous):

oooooh thank you i get it now